# Reassessing Sub-Neptune Structure, Radii, and Thermal Evolution

### v1

@@ 3.5.4 Consequence on the Radius Calculation
To provide a quantitative view of how different assumptions impact the planetary radius, we summarize the relative difference between models as a function planetary mass and incident bolometric flux in Figure 13. The H/He mass fraction is chosen to be the maximum amount allowed from the boil-off mass loss, as in Table 5. Note that our results set the largest possible relative difference at a certain mass and bolometric flux, as a planet born with an initial H/He mass fraction smaller than the value in Table 5 is not vulnerable to boil-off and less sensitive to the variable gravity effect. Our standard model uses the Guillot (2010) T-P profile and includes the variable gravity effect. We compare the effects of variable gravity versus constant gravity, as well as an isothermal atmosphere versus the Guillot atmosphere. The relative differences are calculated at 10 Myr. However, we find that the relative differences vary only by a factor of order unity if we evolve the models to 1 Gyr.
For the 20 mbar radius (top panel), we find that the variable gravity effect is more significant for low-mass planets than for heavier planets, leading to a relative difference on the order of 10%. The influence of the choice of the T-P profile is equally significant. For the 1 nbar radius (bottom panel), the variable gravity effect results in a greater relative difference compared to the 20 mbar radius, reaching up to 60%. The choice of T-P profile also plays an important role, except for planets with $10F_{\oplus}$ and high masses because the average temperature of the Guillot profile below 1 nbar is closer to $T_{\rm{eq}}$, which makes the isothermal atmosphere a reasonable approximation for this specific flux value.
@@
Additionally, we present the optical radius at 1 Gyr calculated with the standard model as a function of rock/iron core mass, incident bolometric flux and H/He mass fraction in Table 7. Comprehensive radius grids, including the RCB radius, optical radius and 1 nbar radius for both pure H/He envelopes as well as 50$\times$ solar metallicity cases, across a broader range of planetary ages, are available online  [^1] https://zenodo.org/records/13985500 for further reference.
@@ 3.6 Implications for Inferred H/He Mass Fractions
In Figure 15, we compare two models assessed with identical planetary parameters and initial conditions with our new model (solid) using parameterized convection and the isothermal assumption (dashed). The EOS for the rock/iron core for the isothermal model follows the work by Lopez et al. (2012). The top panel presents the radii evolution at both the RCB in black and the rock/iron core surface (MEB) in gray, while the bottom panel shows the evolution of the surface temperature of the rock/iron core. Note that our model predicts a much hotter central temperature, often several times higher than that of the isothermal model. Initially, our model predicts a larger RCB radius because the core is molten, causing a 3% size difference compared to the isothermal models. The cooler initial core temperature stems from the shallower location of the rock/iron surface. Early in the evolution, the planet in our model remains 10% larger with a hotter interior due to the contribution of rock/iron luminosities. However, this effect diminishes over time as mantle convection becomes less efficient due to solidification. Notably, after 8 Gyr, the planet’s envelope shrinks rapidly when the mantle fully solidifies, consistent with the processes described in Section 3.4.
@@
Note. — We present a comparison of planetary composition predictions between our model (rightmost columns) and the Lopez et al. (2012) model for Kepler-11 planets. The key planetary parameters used for the evaluation, total mass $M_{p}$, optical radius $R_{p}$ and incident bolometric flux $F_{bol}$, are sourced from the TTV measurements by Lissauer et al. (2011). Our model predicts the envelope mass fraction $f_{\rm{env}}$ and the mass fraction of the radiative atmosphere $f_{\rm{atm}}$ independently. These combine to give the total H/He mass fraction $f_{\rm{HHe}}=f_{\rm{env}}+f_{\rm{atm}}$. In most cases, the envelope mass dominates the total H/He mass, except for planets with a very thin envelope, where $f_{\rm{env}}\sim 0.1\%$. The surface pressure $P_{\rm{surf}}$ is evaluated at the RCB for H/He-rich sub-Neptunes and at the mantle surface for envelope-free super-Earths. In the case of super-Earths, we model the structure by assuming a radiative atmosphere on top of the rock/iron cores. This treatment allows us to extend our approach to these envelope-free planets, offering a more refined prediction for their surface conditions and interior structures.
@@ 4.1 Enhanced Boil-Off Mass Loss
However, the influence of thermal inflation and rock/iron luminosity is considered limited due to several reasons: (1) Since most of the initial H/He mass is lost for low-mass planets, a factor of 3 enhancement in the final mass fraction corresponds to a minimal change of less than 10% in the time-integrated mass loss. (2) The reduction in final mass fraction due to rock/iron luminosity is balanced by an increase in envelope specific entropy, leading to a small difference (less than 10%) in the post-boil-off radius. (3) The final mass fraction is determined by the specific entropy at the end of boil-off. After inflation, the increased radiative luminosity from the inflated interior accelerates cooling, effectively shutting off the boil-off. As a result, the final specific entropy is only slightly higher than the initial entropy, by less than $0.5,k_{b}/\rm{atom}$. This leads to a limited enhancement of time-integrated mass loss. (4) Thermal inflation typically occurs late, usually after more than half of the initial H/He mass is lost. Its effects on mass loss become significant only after about 90% of the initial H/He mass is lost (around 1 Myr, marked with cross shapes in Figure 16). This timing aligns with the expectation that heating from the rock/iron interior becomes influential on the thermal evolution of the envelope once the thermal energy of the rock/iron exceeds the gravitational binding energy of the envelope, which correspond to an inconsiderable envelope mass. (5)While the energy-limited wind in Eq. 14 exhibits stronger coupling with thermal inflation compared to a non-energy-limited Parker wind, the coupling is still relatively weak, as the mass loss occurs about one order of magnitude faster than the thermal evolution timescale throughout most of the boil-off phase. Strong coupling only happens when the thermal inflation timescale is shorter than both the mass loss timescale $t_{\dot{M}}$ and the planetary age $t$, defined as:
@@ 4.2 Presence of Dynamo in the Liquid Iron
We employ this form for $U$, noting that variations in scaling laws typically affect $R_{m}$ only by an order of unity. Here, $L$ is the length scale, taken to be the core radius $R_{c}$, and $\Omega$ is the rotational rate, assumed to correspond to a rotation period of 1 day. The precise value of $\Omega$ does not significantly affect our results. The temperature scale height $H_{T}=c_{P}/(\alpha g)$ is evaluated at each mass shell, with the average value taken for the analysis. The convective energy flux $F_{\rm{conv}}$ that triggers dynamo operation only occurs when the conductive flux is insufficient to transport the cooling energy of the core. This leads to:
@@
In Figure 17, we illustrate the evolution of $R_{m}$ on the left and $F_{\rm{conv}}$ on the right. Each panel varies rock/iron mass, H/He mass fraction, and incident bolometric flux, with different parameters indicated by distinct colors. A wide range of values for $k_{c}$ has been reported from simulation studies. Here, we evaluate an intermediate value of $4\times 10^{6}\ \rm{erg\ cm^{-1}\ K^{-1}}$ (solid) and a high value of $1\times 10^{7}\ \rm{erg\ cm^{-1}\ K^{-1}}$ (dashed) to assess their effects. There are two key limits that constrain $R_{m}$: if the iron core is completely solidified, dynamo action ceases. Additionally, the liquid outer core may stop being convective when conduction dominates energy transport, at which point dynamo action would also cease.
Order of magnitude calculations indicate that as long as a convective layer is present in the core, $R_{m}$ can readily exceed the critical value. Our numerical results show that $R_{m}$ typically ranges from $10^{3}-10^{5}$, well surpassing the critical threshold before $F_{\rm{conv}}$ drops below $10^{-3}\ \rm{erg\ cm^{-2}}$. This suggests that none of the modeled planets are limited by core solidification, due to the long-lasting liquid phase in the iron core. With the intermediate thermal conductivity $k_{c}$ (solid), the convective flux is always sufficient to sustain a dynamo for all modeled planets. However, the planets with the lowest mass (1 $M_{\oplus}$, top) and the smallest H/He mass (0.1%, middle) become limited by convective flux after a few Gyr. After this point, $R_{m}$ can become sensitive to $k_{c}$. As $k_{c}$ increases (dashed), core convection completely ceases.
After the transition, the strength of iron convection sharply declines in the presence of efficient conduction. This transitions coincides with the complete solidification of the mantle surface (see right panels of Figure 7), as the slowed mantle cooling correspondingly decelerate the core evolution, reducing the CMB flux and quenching core convection. Before the mantle surface solidifies, the conductive flux in the core remains negligible compared to the convective flux.
@@ 4.3 Other Complications and Uncertainties
As any modeling framework evolves towards a closer approximation of reality, some aspects can receive detailed attention, and other aspects are left to later work. Here, we did not consider the complex interaction between the magma ocean and the overlying envelope, which can dynamically modify both their compositions and temperature profiles. These processes, including volatile exchange at the MEB and the oxidation of hydrogen in the envelope that leads to water contamination, have been explored in numerous studies. The evolution of water-rich planets demonstrates different observable features compared to that of water-free sub-Neptunes (Zeng et al., 2019; Luque & Pallé, 2022). Such effects influence gas retention in the envelope, with feedback potentially impacting thermal evolution and solidification of the rock/iron core, particularly in planets with thin envelopes. However, it remains unclear whether the water dissolves into the iron core (Luo et al., 2024), the mantle (Schlichting & Young, 2022; Kite et al., 2020; Vazan et al., 2022), or the envelope, or forms a distinct water layer (Rogers & Seager, 2010). These uncertainties highlight future research directions.
@@
Further uncertainty in radius evolution arises from the potential condensation of silicate vapor in the envelope, sourced from pebble accretion. Over Gyr timescales, silicate vapor may settle into the rock/iron core, releasing energy and potentially inflating the planetary radius (Vazan & Ormel, 2023). This process could decrease the estimated H/He mass fraction and enhance the mass loss rate (Vazan et al., 2024).
@@ 5 Conclusions
    In low-mass planets, mantle solidification occurs from the bottom up, whereas in heavier planets, it initiates in the lower middle, with the phase change interfaces expanding quickly downward and more slowly upward. This behavior may be significantly influenced by the EOS employed, warranting further studies for confirmation.

### v2

@@ 3.5.4 Consequence on the Radius Calculation
To provide a quantitative view of how different assumptions impact the planetary radius, we summarize the relative difference between models as a function rock/iron mass and incident bolometric flux in Figure 15. The H/He mass fraction is chosen to be the maximum amount allowed from the boil-off mass loss, as in Table 5. Note that our results set the largest possible relative difference at a certain mass and bolometric flux, as a planet born with an initial H/He mass fraction smaller than the value in Table 5 is not vulnerable to boil-off and less sensitive to the variable gravity effect. Our standard model uses the Guillot (2010) T-P profile and includes the variable gravity effect. We compare the effects of variable gravity versus constant gravity, as well as an isothermal atmosphere versus the Guillot atmosphere. The relative differences are calculated at 10 Myr. However, we find that the relative differences vary only by a factor of order unity if we evolve the models to 1 Gyr.
For the 20 mbar radius (top panel), we find that the variable gravity effect is more significant for low-mass planets than for heavier planets, leading to a relative difference on the order of 10%. The influence of the choice of the T-P profile is equally significant. For the 1 nbar radius (bottom panel), the variable gravity effect results in a greater relative difference compared to the 20 mbar radius, reaching up to 60%. The choice of T-P profile also plays an important role, except for planets with $10F_{\oplus}$ (second dashed curve from the left in the bottom panel) because the average temperature of the Guillot profile below 1 nbar is closer to $T_{\rm{eq}}$ than in other models (as shown in Figure 2, where the average temperature is hotter than $T_{\rm{eq}}$), which makes the isothermal atmosphere a reasonable approximation for this specific flux value. For this curve, the 1 nbar radius is slightly overestimated relative to our standard atmosphere model, resulting in an opposite slope compared to other curves. In all other cases, we find that these physical choices systematically underestimate atmospheric radii due to the underestimated temperature and overestimated gravity.
@@
Additionally, we present the optical radius at 1 Gyr calculated with the standard model as a function of rock/iron core mass, incident bolometric flux and H/He mass fraction in Table 7. Comprehensive radius grids, including the RCB radius, optical radius and 1 nbar radius for both pure H/He envelopes as well as 50$\times$ solar metallicity cases, across a broader range of planetary ages, are available online  [^1] https://doi.org/10.5281/zenodo.13985499 for further reference. In the 1 nbar and optical radius tables, planets susceptible to boil-off are excluded, as hydrostatic equilibrium no longer holds in the radiative atmosphere. A visualization tool is also available for interpolating the presented data (Aguichine, 2024).
@@ 3.6 Implications for Inferred H/He Mass Fractions
In Figure 17, we compare two models assessed with identical planetary parameters and initial conditions with our new model (solid) using parameterized convection and the isothermal assumption (dashed). The EOS for the rock/iron core for the isothermal model follows the work by Lopez et al. (2012). The top panel presents the radii evolution at both the RCB in black and the rock/iron core surface (MEB) in gray, while the bottom panel shows the evolution of the surface temperature of the rock/iron core. Note that our model predicts a much hotter central temperature, often several times higher than that of the isothermal model. Initially, our model predicts a larger RCB radius because the rock/iron core is molten, causing a 3% size difference compared to the isothermal models. The cooler initial core temperature stems from the shallower location of the rock/iron surface. Early in the evolution, the planet in our model remains 10% larger with a hotter interior due to the contribution of rock/iron luminosities. However, this effect diminishes over time as mantle convection becomes less efficient due to solidification. Notably, after 8 Gyr, the planet’s envelope shrinks rapidly when the mantle fully solidifies, consistent with the processes described in Section 3.4.
@@
Note. — We present a comparison of planetary composition predictions between our model (rightmost columns) and the Lopez et al. (2012) model for Kepler-11 planets. The key planetary parameters used for the evaluation, total mass $M_{p}$, optical radius $R_{p}$ and incident bolometric flux $F_{rm{bol}}$, are sourced from the TTV measurements by Lissauer et al. (2011). Our model predicts the envelope mass fraction $f_{\rm{env}}$ and the mass fraction of the radiative atmosphere $f_{\rm{atm}}$ independently. These combine to give the total H/He mass fraction $f_{\rm{HHe}}=f_{\rm{env}}+f_{\rm{atm}}$. In most cases, the envelope mass dominates the total H/He mass, except for planets with a very thin envelope, where $f_{\rm{env}}\sim 0.1\%$. The surface pressure $P_{\rm{surf}}$ is evaluated at the RCB for H/He-rich sub-Neptunes and at the mantle surface for envelope-free super-Earths. In the case of super-Earths, we model the structure by assuming a radiative atmosphere on top of the rock/iron cores. This treatment allows us to extend our approach to these envelope-free planets, offering a more refined prediction for their surface conditions and interior structures.
@@ 4.1 Enhanced Boil-Off Mass Loss
However, the influence of thermal inflation and rock/iron luminosity is considered limited due to several reasons: (1) Since most of the initial H/He mass is lost for low-mass planets, a factor of 3 enhancement in the final mass fraction corresponds to a minimal change of less than 10% in the time-integrated mass loss. (2) In many cases, boil-off alone is sufficient to completely strip a sub-Neptune’s envelope, even without considering the contribution from core luminosity. (3) The reduction in final mass fraction due to rock/iron luminosity is balanced by an increase in envelope specific entropy, leading to a small difference (less than 10%) in the post-boil-off radius. (4) The final mass fraction is determined by the specific entropy at the end of boil-off. After inflation, the increased radiative luminosity from the inflated interior accelerates cooling, effectively shutting off the boil-off. As a result, the final specific entropy is only slightly higher than the initial entropy, by less than $0.5,k_{b}/\rm{atom}$. This leads to a limited enhancement of time-integrated mass loss. (5) Thermal inflation typically occurs late, usually after more than half of the initial H/He mass is lost. Its effects on mass loss become significant only after about 90% of the initial H/He mass is lost (around 1 Myr, marked with cross shapes in Figure 18). This timing aligns with the expectation that heating from the rock/iron interior becomes influential on the thermal evolution of the envelope once the thermal energy of the rock/iron exceeds the gravitational binding energy of the envelope, which correspond to an inconsiderable envelope mass. (6) While the energy-limited wind in Eq. 14 exhibits stronger coupling with thermal inflation compared to a non-energy-limited Parker wind, the coupling is still relatively weak, as the mass loss occurs about one order of magnitude faster than the thermal evolution timescale throughout most of the boil-off phase. Strong coupling only happens when the thermal inflation timescale is shorter than both the mass loss timescale $t_{\dot{M}}$ and the planetary age $t$, defined as:
@@ 4.2 Presence of Dynamo in the Liquid Iron
We employ this form for $U$, noting that variations in scaling laws typically affect $R_{m}$ only by an order of unity. Here, $L$ is the length scale, taken to be the core radius $R_{c}$, and $\Omega$ is the rotational rate, assumed to correspond to a rotation period of 1 day. The precise value of $\Omega$ does not significantly affect our results. The temperature scale height $H_{T}=c_{P}/(\alpha g)$ is evaluated at each mass shell, with the average value taken for the analysis. The convective energy flux $F_{\rm{conv}}$ that triggers dynamo operation only occurs when the conductive flux along the adiabat is insufficient to transport the cooling energy of the core. This leads to:
@@
In Figure 19, we illustrate the evolution of $R_{m}$ on the left and $F_{\rm{conv}}$ on the right. Each panel varies rock/iron mass, H/He mass fraction, and incident bolometric flux, with different parameters indicated by distinct colors. A wide range of values for $k_{c}$ has been reported from simulation studies. Here, we evaluate an intermediate value of $4\times 10^{6}\ \rm{erg\ s^{-1}\ cm^{-1}\ K^{-1}}$ (Konôpková et al., 2016, solid) and a high value of $1\times 10^{7}\ \rm{erg\ s^{-1}\ cm^{-1}\ K^{-1}}$ (Pozzo et al., 2012, dashed) to assess their effects. There are two key limits that constrain $R_{m}$: if the iron core is completely solidified, dynamo action ceases. Additionally, the liquid outer core may stop being convective when conduction dominates energy transport, at which point dynamo action would also cease.
All modeled planets in the figure retain a liquid outer iron core by 10 Gyr. Moreover, order of magnitude calculations indicate that as long as a convective layer is present in the liquid iron, $R_{m}$ can readily exceed the critical value. Our numerical results show that $R_{m}$ typically ranges from $10^{3}-10^{5}$, well surpassing the critical threshold before $F_{\rm{conv}}$ drops below $10^{-3}\ \rm{erg\ cm^{-2}}$.
However, we find that the presence of convection in the liquid-phase iron core is highly sensitive to thermal conductivity. With the intermediate thermal conductivity $k_{c}$ (solid), the convective flux is always present to sustain a dynamo for all modeled planets. As $k_{c}$ increases (dashed), iron convection ceases at a transition time of a few Gyr for the lowest-mass planets (1 $M_{\oplus}$, top) and those with the smallest H/He mass fraction (0.1%, middle). This transition coincides with the complete solidification of the mantle surface (see right panels of Figure 8), as the slowed mantle cooling consequently decelerates the core’s thermal evolution, reducing the CMB flux and ultimately quenching core convection. Notably, at both $k_{c}$ values, the conductive flux within the iron core remains negligible compared to the convective flux until the mantle surface solidifies.
In summary, we suggest that constraining the thermal conductivity of the iron core is crucial for improving our understanding of dynamo operation in low-mass, thin-envelope planets.
@@ 4.3.1 Interaction between species
As any modeling framework evolves towards a closer approximation of reality, some aspects can receive detailed attention, and other aspects are left to later work. Here, we did not consider the complex interaction between the magma ocean and the overlying envelope, which can dynamically modify both their compositions and temperature profiles. These processes, including volatile exchange at the MEB and the oxidation of hydrogen in the envelope that leads to water contamination, have been explored in numerous studies. The evolution of water-rich planets demonstrates different observable features compared to that of water-free sub-Neptunes (Zeng et al., 2019; Luque & Pallé, 2022). Such effects influence gas retention in the envelope, with feedback potentially impacting thermal evolution and solidification of the rock/iron core, particularly in planets with thin envelopes. However, it remains unclear whether the water dissolves into the iron core (Luo et al., 2024), the mantle (Schlichting & Young, 2022; Kite et al., 2020; Vazan et al., 2022), or the envelope, or forms a distinct water layer (Rogers & Seager, 2010). Such volatile impurities in the rock/iron core may alter its thermodynamic properties (e.g., melting temperature, heat capacity, and thermal conductivity), introducing uncertainties in the planet’s thermal evolution. These complexities highlight future research directions.
4.3.2 Differentiation
At early ages during the formation phase, a sub-Neptune’s rock/iron core is likely in a homogeneous mixture state, owing to frequent impacts during accretion. Over time, the iron content from impactors separates from the silicate material due to the density contrast, eventually forming a distinct iron core beneath a silicate mantle. This redistribution of mass releases energy comparable in magnitude to the total thermal energy of the rock/iron component $E_{\rm{ric}}$. However, core-mantle differentiation is not explicitly modeled in this work, nor in most previous sub-Neptune evolution models.
Although extensively studied in terrestrial planets, the differentiation process remains poorly understood in the context of sub-Neptunes. In this context, we consider two possible scenarios based on current theoretical understanding.
In the classical Earth-like models of differentiation, iron droplets rain out within the surface magma ocean, accumulating to form a liquid iron pond, and subsequently, sink through the overlaying solid mantle into the developing iron core (Stevenson, 1990; Rubie et al., 2007). However, our results suggest that sub-Neptunes typically retain a fully molten mantle for Gyr timescales (Figure 8), which implies that iron core formation may occur rapidly and directly, within a nearly instantaneous timescale. In this scenario, the energy released by differentiation contributes only modestly to the initial entropy of the envelope and interior, since the envelope energy $E_{\rm{env}}$ usually dominates over the energy of the rock/iron core $E_{\rm{ric}}$. We tested this effect by shortening the initial contraction timescale to 3 Myr, which results in a 10–20% reduction in the post-boil-off envelope mass fraction. This alters the evolution of the envelope’s specific entropy by only a negligible amount (see middle-left panel of Figure 4).
In contrast, Wahl & Militzer (2015) propose a different differentiation mechanism based on the miscibility of iron and silicate/oxide mantle. They identify a critical temperature above which a homogeneous mixture of iron and mantle materials is stable. This critical temperature is relatively low and is reached within the solid mantle phase at high pressures, implying that differentiation could be delayed until later evolutionary stages. If this delayed scenario occurs, the differentiation energy would be released on Gyr timescales, when $E_{\rm{ric}}$ may exceed $E_{\rm{env}}$, potentially leading to elevated interior temperatures and delayed cooling. Given these uncertainties, the role of delayed differentiation in sub-Neptunes warrants further investigation in future studies.
4.3.3 Silicate vapor
@@
Further uncertainty in radius evolution arises from the potential condensation of silicate vapor in the envelope, sourced from planetary accretion. Over Gyr timescales, silicate vapor may settle into the rock/iron core, releasing energy and potentially inflating the planetary radius (Vazan & Ormel, 2023). This process could decrease the estimated H/He mass fraction and enhance the mass loss rate (Vazan et al., 2024).
@@ 5 Conclusions
    In low-mass planets, mantle solidification progresses strictly from the bottom up. In contrast, in heavier planets, it initiates in the middle of the lower mantle and quickly dominates the entire lower mantle. After this point, the lower mantle solidifies nearly simultaneously and the phase change interface ($\phi=0.5$) advances upward. This behavior may be significantly influenced by the EOS employed, warranting further studies for confirmation.